OpenStudy (tester97):

Use the Remainder Theorem to find the remainder when P(x) = x4 – 9x3 – 5x2 – 3x + 4 is divided by x + 3. @wolf1728 @Luigi0210

OpenStudy (luigi0210):

I'll let @wolf1728 get this one, good luck.

OpenStudy (wolf1728):

Wow - I was using the Internet to find Remainder Theorem - Wikipedia has it - I'm totally unfamiliar with it.

OpenStudy (tester97):

NOOOO D;

OpenStudy (wolf1728):

Well maybe I can study it a little bit more.

OpenStudy (luigi0210):

I'll just bump in to get you guys started. I think you can use the remainder theorem with synthetic division too, so start with that.

OpenStudy (tester97):

wait actually i think i got it is this correct? because if you use remainder theorem you use p(-3) right? if so then this is what i came up with? P(-3) = (-3)^4 - 9(-3)^3 - 5(-3)^2 - 3(-3)+4 P(-3) = 81 - 9(-27) - 5(9) + 9 + 4 P(-3) = 81 + 243 - 45 + 9 + 4 P(-3) = 324 - 45 + 13 P(-3) = 337 - 45 P(-3) = 292 Then divide 292 by x+3?

OpenStudy (wolf1728):

1 -9 -5 -3 4 -1 -3 That's as far as I got with synthetic division.

OpenStudy (tester97):

oh ok thanks for trying though :)

OpenStudy (wolf1728):

Well Luigi is still here.

OpenStudy (luigi0210):

Here's how it would look like using synthetic:  But I feel like I messed up somewhere..